The side of square are initially 4 inches in length and increase at a constant rate of 3 inches per secondOne corner of suqare moves along a line L at a speed of 2 inches per second for 5 second.During this time the square makes one revolution about p revolving at a constant rate and remaining always perpendicular to L. explain how carefully how to compute volume of solid generated by square.Carry out the computation you describe


1
Expert's answer
2021-06-08T18:36:01-0400

The side of the square moves at a constant rate of 3 inches per second.

initially the side of the square 4 inches after 5seconds it has 4 + 3 * 5 = 19.

The corner of the square moves along a line L at a speed of 2 inches per second after 5

seconds it has 10 inches. L is perpendicular to square. So L is the height of solid.

V = 19 * 19 * 10 = 3610